# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ 2.3.1 Optical Character Recognition (OCR)
PDF is the predominant format for disseminating financial content such as reports, regulatory filings, and legal documents. However, current financial vision tasks xie2024open largely focus on image understanding rather than the complex structure of document-based financial information. To address this gap, we introduce the first financial OCR task aimed at converting scanned financial PDF files (in image format) into structured HTML text. We further construct two novel datasets in English and Spanish, the two most widely used languages in financial communication within the United States.
Task definition
@@
The OCR task is defined as a structured information extraction problem from document images. Each financial PDF document is segmented into a set of page-level images $\{I_{1},I_{2},\ldots,I_{n}\}$, where each image $I_{i}$ corresponds to a single page. The model processes each image individually and generates a corresponding HTML-formatted text sequence $T_{i}$, such that $T_{i}=\text{OCR}(I_{i})$. The objective is not only to recognize the textual content but also to recover the structural and semantic layout embedded in the document, including headings, tables, figures, and paragraph segmentation.
Data source
Following the task design, we construct two novel datasets: EnglishOCR and SpanishOCR. The EnglishOCR dataset was built using U.S. SEC EDGAR filings, which are primarily available in HTML format. For documents without a native PDF version, we used wkhtmltopdf to render PDF files from the corresponding HTML. Each document was then parsed into page-level PNG images along with its corresponding HTML content. To align each image with the most relevant HTML snippet, we applied cosine similarity using a Sentence-BERT (SBERT) model on OCR-extracted text and HTML sentences.
The SpanishOCR dataset was constructed using source PDFs obtained from Peruvian public regulatory documents. Each document was parsed into page-level PNG images and the corresponding HTML content was generated by applying OCR to extract text and then wrapping it in appropriate HTML tags to preserve basic document structure.
Evaluation metric
We evaluate model outputs using ROUGE-1 lin-2004-rouge , a unigram overlap metric that measures lexical similarity between model predictions and references.
2.4 Audio Benchmark
To evaluate LLM and AudioLLM’s performance on financial audio scenarios, we include two English audio datasets from FinAudio cao2025finaudiobenchmarkaudiolarge under the TG task, covering Automatic Speech Recognition (ASR) and speech summarization tasks. For ASR, the MDRM qin2019you dataset consists of short financial audio clips from earnings conference call recordings, comprising 22,208 clips (87 hours). This dataset assesses transcription accuracy using Word Error Rate (WER) park2008empirical . For Speech Summarization, the FinAudioSum cao2025finaudiobenchmarkaudiolarge dataset includes 64 long financial audio recordings (55 hours) paired with abstractive summaries. Performance is evaluated using ROUGE-L lin-2004-rouge . More details in Appendix G.
3 Experiments
Access Modality Language Domain Target Language Model MOF Class Closed Multimodal Multilingual General English GPT-4o hurst2024gpt Class III - Open Model - In progress (17%) Text Multilingual General English GPT-o3-mini openaio3mini Class III - Open Model - In progress (0%) Open Multimodal Multilingual General English meta-llama/Llama-4-Scout-17B-16E-Instruct llama4meta Class III - Open Model - In progress (0%) General English google/gemma-3-4b-it gemmateam2025gemma3technicalreport Class III - Open Model - In progress (17%) General English google/gemma-3-27b-it gemmateam2025gemma3technicalreport Class III - Open Model - In progress (17%) General English Qwen/Qwen2.5-Omni-7B Qwen2.5-Omni Class III - Open Model - In progress (67%) Text Multilingual General English meta-llama/Llama-3.1-70B-Instruct dubey2024llama Class III - Open Model - In progress (0%) General English Deepseek-V3 liu2024deepseek Class III - Open Model - In progress (17%) General English Qwen/Qwen2.5-32B-Instruct qwen2.5 Class III - Open Model - In progress (33%) General Japanese cyberagent/DeepSeek-R1-Distill-Qwen-32B-Japanese cyberagent-deepseek-r1-distill-qwen-32b-japanese Class III - Open Model - In progress (50%)1 Financial Chinese Duxiaoman-DI/Llama3.1-XuanYuan-FinX1-Preview xuanyuan2024 Class III - Open Model - In progress (0%)1 Financial Spanish TheFinAI/FinMA-ES-Bilingual 10.1145/3637528.3671554 Class III - Open Model - In progress (50%)1 Financial Greek TheFinAI/plutus-8B-instruct peng2025plutus Class III - Open Model - In progress (17%)1 English Financial English TheFinAI/finma-7b-full xie2023pixiu Class III - Open Model - In progress (67%)1 Vision Multilingual General English Qwen/Qwen-VL-Max Qwen-VL Class III - Open Model - In progress (0%) General English Deepseek-VL-7B-Chat lu2024deepseekvl Class III - Open Model - In progress (17%) English General English llava-hf/llava-v1.6-vicuna-13b-hf liu2023improved Class III - Open Model - Qualified & Class II - Open Tooling - In progress (80%)1 Audio Multilingual General English Whisper-V3 pmlr-v202-radford23a Class III - Open Model - In progress (67%) General English Qwen2-Audio-7B chu2024qwen2audiotechnicalreport Class III - Open Model - In progress (67%) General English Qwen2-Audio-7B-Instruct chu2024qwen2audiotechnicalreport Class III - Open Model - In progress (67%) General English SALMONN-7B tang2024salmonngenerichearingabilities Class III - Open Model - In progress (50%)1 General English SALMONN-13B tang2024salmonngenerichearingabilities Class III - Open Model - In progress (50%)1
  • 1
Models
We include 22 models (Table 3) across text, vision, audio, and multimodal tasks, considering their openness under the MOF framework mof2024white  [^5] Please check Appendix I for more details. For comparison, we include closed-source models, GPT-4o hurst2024gpt and GPT-o3-mini openaio3mini , which do not meet MOF Class III. The open-source models mostly fall under Class III, including multilingual multimodal models (Llama-4 llama4meta , Gemma-3 series gemmateam2025gemma3technicalreport , Qwen2.5-Omni Qwen2.5-Omni ), text-only multilingual models (Llama-3.1-70B dubey2024llama , Deepseek-V3 liu2024deepseek , Qwen2.5 qwen2.5 ), and monolingual financial models (FinMA xie2023pixiu , XuanYuan xuanyuan2024 , DeepSeek-R1-Distill-Japanese cyberagent-deepseek-r1-distill-qwen-32b-japanese , FinMA-ES 10.1145/3637528.3671554 , Plutus peng2025plutus ). For vision-language, we assess Qwen-VL-Max Qwen-VL , DeepSeek-VL lu2024deepseekvl , and LLaVA-v1.6-Vicuna-13B liu2023improved , the only model partially meeting Class II. Audio-language models include Whisper-V3 pmlr-v202-radford23a , Qwen2-Audio-7B, Qwen2-Audio-7B-Instruct chu2024qwen2audiotechnicalreport , and SALMONN-7B/13B tang2024salmonngenerichearingabilities .
Implementation details
@@ 3 Experiments
To ensure evaluation integrity, we customize our evaluation pipeline based on the LM Evaluation Harness eval-harness . GPT and Together-hosted models (DeepSeek, Llama-4, Gemma-3) are queried via official APIs with temperature set to 0. All other open-source models are deployed and evaluated locally using vLLM kwon2023efficient on GPUs. Including the OpenAI and Together AI API costs, the total expenditure amounts to approximately \$80,000.
4 Results
Table 4: Standardized performance of evaluated LLMs on the MultiFinBen.
Modality Language Task Dataset GPT-4o o3-Mini Deepseek-V3 Llama-4-Scout Llama-3.1-70B Gemma-4B Gemma-27B Qwen2.5-32B Qwen2.5-Omni FinMA-7B XuanYuan R1-Qwen32B-JA FinMA-ES Plutus-8B Qwen-VL-MAX LLaVA-1.6 Deepseek-VL Whisper-V3 Qwen2-Audio Qwen2-Audio-Inst SALMONN-7B SALMONN-13B Text EN IE SC 88.00 0.00 0.00 20.73 87.00 0.69 0.00 22.28 18.61 56.62 24.09 15.92 52.70 19.79 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 IE FinRED 3.00 0.00 0.00 0.00 0.00 0.00 0.00 0.37 0.09 0.00 0.00 0.00 0.00 0.75 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 IE FINER-ORD 78.00 9.58 0.18 2.23 18.00 0.00 0.00 28.30 8.30 0.04 0.00 0.00 0.00 5.35 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA Headlines 78.00 0.00 47.32 84.33 60.00 0.00 0.00 85.42 82.23 97.08 85.10 82.21 94.69 71.14 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA TSA 61.00 0.00 0.85 63.82 42.00 32.34 32.34 42.98 43.40 81.70 85.11 60.00 86.38 54.89 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA XBRL-Math 68.00 68.89 67.78 27.78 62.00 0.00 11.11 64.44 44.44 7.78 3.33 2.22 6.67 14.44 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA FinQA 5.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 7.41 0.00 0.00 0.00 1.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA TATQA 0.00 0.00 0.00 0.36 44.00 0.00 0.00 0.05 1.73 4.14 0.00 0.00 0.00 15.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG ECTSUM 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG EDTSUM 25.00 19.13 16.80 16.59 18.00 0.98 0.10 20.16 23.89 19.92 12.49 8.06 2.06 13.61 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 RM CCF 52.50 50.00 50.62 51.34 50.00 50.93 50.00 52.94 50.31 50.05 50.00 50.00 51.18 50.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 PO BigData22 48.50 50.00 50.93 46.91 50.00 50.75 50.00 49.89 51.82 50.80 53.12 50.00 52.12 50.26 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 DM MSFT 41.32 65.06 0.00 0.00 72.25 74.03 79.97 49.32 0.00 0.00 68.81 74.50 66.53 65.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 42.18 20.20 18.04 24.16 38.71 16.13 17.19 32.01 24.99 28.89 29.39 26.38 31.72 27.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 ZH IE RRE 63.25 0.00 67.52 54.70 46.15 36.75 36.75 8.55 7.69 0.85 2.56 0.85 0.85 2.56 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA AIE 82.26 0.00 82.01 80.99 76.80 33.82 33.82 83.03 80.17 40.81 10.04 4.32 21.55 54.48 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA LNE 63.30 0.00 58.72 55.50 41.28 9.17 9.17 57.80 59.17 29.82 22.48 12.84 32.11 26.61 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA FinancialIQ 32.53 0.00 35.52 66.83 62.71 25.19 25.20 77.09 65.32 26.21 57.07 34.70 31.48 40.52 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 60.34 0.00 60.94 64.51 56.74 26.23 26.24 56.62 53.09 24.42 23.04 13.18 21.50 31.04 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 JA TA chabsa 0.00 0.00 0.00 48.43 32.17 8.98 23.96 4.54 44.35 46.94 47.59 23.96 57.36 34.62 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 ES TA MultiFin 61.74 0.00 53.91 62.17 48.26 22.17 22.17 46.52 46.96 43.04 31.74 12.61 44.78 51.30 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA TSA 0.39 0.00 29.17 52.29 24.29 63.04 63.46 31.63 46.46 31.03 68.19 63.38 16.64 51.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA EFPA 31.14 0.00 18.86 67.54 66.67 25.44 25.44 65.79 55.70 32.46 65.79 25.44 91.67 48.25 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG FNS-2023 25.94 18.11 0.00 9.61 12.14 0.00 0.00 5.93 7.50 1.64 5.71 10.62 1.65 9.27 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 29.80 4.53 25.49 47.90 37.84 27.66 27.77 37.47 39.16 27.04 42.86 28.01 38.69 40.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 EL IE GRFinNUM 9.18 20.98 7.43 49.12 46.34 0.00 0.00 36.77 0.40 0.00 0.00 0.00 0.00 70.06 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA GRFinQA 78.22 0.00 50.00 74.22 64.44 22.67 22.67 60.44 48.89 25.33 57.78 28.44 23.11 64.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG GRFNS-2023 25.50 16.95 37.72 16.90 13.61 0.24 0.21 9.71 5.60 11.20 6.48 14.45 3.56 34.46 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA GRMultiFin 59.26 0.00 61.11 55.56 50.00 38.89 38.89 70.37 38.89 35.19 53.70 40.74 35.19 72.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 43.04 9.48 39.07 48.95 43.60 15.45 15.44 44.32 23.45 17.93 29.49 20.91 15.47 60.19 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 BI TG DOLFIN 92.29 90.13 86.26 89.17 92.13 35.92 35.92 92.29 91.80 69.24 91.60 71.81 66.57 91.59 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 MU QA PolyFiQA-Easy 9.15 10.54 42.58 14.80 26.06 0.00 0.00 19.77 18.12 2.14 2.73 12.05 0.70 6.46 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA PolyFiQA-Expert 3.90 5.05 31.40 12.23 17.88 0.00 0.00 17.18 14.45 4.05 0.78 8.45 0.00 8.02 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 6.53 7.80 36.99 13.52 21.97 0.00 0.00 18.48 16.29 3.10 1.76 10.25 0.35 7.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 40.91 14.64 30.92 38.76 42.21 18.34 19.35 38.05 32.98 26.74 31.25 24.40 28.95 35.45 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Vision EN IE EnglishOCR 21.38 0.00 0.00 12.39 0.00 10.70 11.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA TableBench 66.70 0.00 0.00 32.30 0.00 28.60 60.90 0.00 74.90 0.00 0.00 0.00 0.00 0.00 55.40 59.30 57.30 0.00 0.00 0.00 0.00 0.00 ES IE SpanishOCR 78.55 0.00 0.00 4.12 0.00 5.60 4.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 55.54 0.00 0.00 16.27 0.00 14.97 25.57 0.00 24.97 0.00 0.00 0.00 0.00 0.00 18.47 19.77 19.10 0.00 0.00 0.00 0.00 0.00 Audio EN TG MDRM-test 97.92 0.00 0.00 0.00 0.00 0.00 0.00 0.00 96.43 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 97.86 96.03 95.32 48.48 49.17 TG FinAudioSum 13.20 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.30 0.00 4.80 0.00 0.00 Average 55.56 0.00 0.00 0.00 0.00 0.00 0.00 0.00 48.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 51.58 48.02 50.06 24.24 24.59 Average 50.67 4.88 10.31 18.34 14.07 11.10 14.97 12.68 35.39 8.91 10.42 8.13 9.65 11.82 6.16 6.59 6.37 17.19 16.01 16.69 8.08 8.20
@@ 4 Results
Table 4  [^6] Results are also visualized on our leaderboard. For more details, refer to Appendix J. https://huggingface.co/spaces/TheFinAI/Open-FinLLM-Leaderboard presents model performance on the MultiFinBen benchmark. For fair comparison, we additionally report modality-balanced overall scores for each model. We highlight several key findings below:
MultiFinBen presents a substantial challenge for state-of-the-art LLMs.
Despite its state-of-the-art performance, GPT-4o achieves an overall score of only 50.67% on the proposed MultiFinBen benchmark, revealing significant limitations even for leading models. The runner-up, Qwen-2.5-Omni-7B, follows with a score of 35.39%, highlighting the steep challenge posed by this comprehensive evaluation. Notably, both models are multimodal and multilingual, underscoring the necessity of such capabilities for competitive performance. In stark contrast, monomodal and monolingual models exhibit markedly inferior results. The best-performing text-only model, Llama-3.1-70B, achieves merely 14.07%. Modality-specific models such as Whisper-V3 (audio-only, 17.19%), Deepseek-VL-7B-Chat (vision-only, 6.37%), and LLaVA-v1.6-Vicuna-13B (English-only, 6.59%) further emphasize the substantial performance gap. These disparities highlight the inherent limitations of models lacking robust cross-modal and cross-lingual reasoning, underscoring the critical need for integrating multilingual and multimodal insights in financial decision-making, as rigorously stress-tested by our MultiFinBen.
Multimodal models show trade-offs across modalities.
while GPT-4o and Qwen-2.5-Omni-7B lead overall on MultiFinBen, their rankings in text-only tasks drop to 2nd and 6th, behind the text-specialized Llama-3.1-70B. This trade-off becomes more pronounced in mid-tier models like gemma-3-4b (11th overall, 13th in text) and gemma-3-27b (7th overall, 12th in text), reflecting the difficulty of maintaining text-specific optimization when broadening to multiple modalities. In contrast, multimodal models decisively outperform unimodal baselines in vision and audio tasks, i.e, GPT-4o achieves 55.54% in vision and 55.56% in audio, far surpassing modality-specific counterparts. The asymmetric performance reveals that while text tasks benefit from data maturity, vision and audio gain more from multimodal integration, making unified models essential for complex financial real-world applications.
(a) Performance across modalities: Audio, Vision, Text.
(b) Performance across languages: EN, ZH, JA, ES, EL, BI, MU.

### v3

@@ 2.4.2 Optical Character Recognition (OCR)
While financial communication heavily relies on PDFs, existing benchmarks Xie et al. (2024b) focus on visual QA or chart reasoning rather than document-structured text recovery. We introduce the first multilingual financial OCR task, targeting end-to-end extraction of structured content from scanned filings, regulatory documents, and financial white papers. This task bridges the gap between visual understanding and textual reasoning, enabling evaluation of multimodal systems on real-world document pipelines.
Task Definition.
@@
The OCR task is defined as a structured information extraction problem from document images. Each financial PDF document is segmented into a set of page-level images $\{I_{1},I_{2},\ldots,I_{n}\}$, where each image $I_{i}$ corresponds to a single page. The model processes each image individually and generates a corresponding HTML-formatted text sequence $T_{i}$, such that $T_{i}=\text{OCR}(I_{i})$, preserving both textual content and document structure (e.g., headings, tables, and paragraphs). The goal is to evaluate structural fidelity and semantic recovery, not just text recognition.
Data Construction.
Following the task design, we construct four novel datasets in a multilingual setting: EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR (Table 2). The EnglishOCR dataset is built using U.S. SEC EDGAR filings [^3] https://www.sec.gov/search-filings, which are primarily distributed in HTML format. For documents without native PDF versions, we use wkhtmltopdf to render PDF files from the corresponding HTML sources. Each PDF is segmented into page-level PNG images aligned with its corresponding HTML content. To match each image with the most relevant HTML snippet, we compute cosine similarity between OCR-extracted text and HTML sentences using a Sentence-BERT (SBERT) model, resulting in 7,961 aligned image–HTML pairs. The JapaneseOCR, SpanishOCR, and GreekOCR datasets are constructed using source PDFs from Japanese Financial Services Agency (FSA) white papers [^4] https://www.fsa.go.jp/en/, Peruvian public regulatory documents [^5] https://www.bvl.com.pe/en/home-general, and Athens Stock Exchange annual company filings [^6] https://www.athexgroup.gr/en/market-data/issuers, respectively. Each PDF is decomposed into page-level PNG images, and the corresponding HTML content is generated by applying OCR to extract content and wrapping it in HTML tags to preserve document structure. We further apply quality-control by removing pages with excessive OCR errors and post-processing malformed outputs, such as repeated or garbled characters. After filtering, these datasets contain 17,586, 12,819, and 6,533 aligned image–HTML pairs, respectively. Detailed statistics are provided in Table 7 and Appendix G.
Evaluation Metric.
We employ ROUGE-1 Lin (2004) to measure lexical and structural overlap between predicted and reference HTML sequences, complemented by human evaluation (Appendix J) for reliable factuality assessment.
3 Experimental Results
To rigorously evaluate the effectiveness of 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: in exposing model limitations and guiding future development, we structure our experiments around four key research questions (RQs): RQ1: How severe are the performance deficits of current LLMs when confronted with the multilingual and multimodal scenarios ubiquitous in real-world finance? RQ2: Are there significant biases or trade-offs in model performance across different modalities? RQ3: Can the model’s monolingual capabilities be effectively generalized to complex multilingual tasks that require cross-lingual comprehensive reasoning? RQ4: How does our difficulty-aware framework help analyze the strengths and weaknesses of the model in more detail?
Evaluation Models.
We evaluate 21 models (Table 4, Appendix C) spanning text, vision, audio, and multimodal modalities, considering their openness under the Model Openness Framework (MOF, Appendix K) White et al. (2024b).
Implementation Details.
@@ 3 Experimental Results
To ensure evaluation integrity and consistency, we customize our evaluation pipeline based on the LM Evaluation Harness Gao et al. (2024). OpenAI and TogetherAI-hosted models (DeepSeek, Llama-4, and Gemma-3) are accessed via their official APIs with temperature set to 0. All other open-source models are deployed and evaluated locally using vLLM Kwon et al. (2023) on GPUs. Including the OpenAI and TogetherAI API costs, the total expenditure amounts to approximately \$80,000.
Key Results.
@@
Table 3 presents model performance on the 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark [^7] Results are also visualized on our leaderboard. For more details, refer to Appendix L.. To ensure fair comparison, we additionally report a modality-balanced overall score for each model. The major findings are summarized as follows.
Modality Language Task Dataset GPT-4o o3-mini Deepseek-V3 Llama-4-Scout Llama-3.1-70B Gemma-4B Gemma-27B Qwen2.5-32B Qwen2.5-Omni FinMA-7B XuanYuan R1-Qwen32B-JA FinMA-ES Plutus-8B LLaVA-1.6 Deepseek-VL Whisper-V3 Qwen2-Audio Qwen2-Audio-Inst SALMONN-7B SALMONN-13B Text EN IE SC 88.00 0.00 0.00 20.73 87.00 0.69 0.00 22.28 18.61 56.62 24.09 15.92 52.70 19.79 0.00 0.00 0.00 0.00 0.00 0.00 0.00 IE FinRED 3.00 0.00 0.00 0.00 0.00 0.00 0.00 0.37 0.09 0.00 0.00 0.00 0.00 0.75 0.00 0.00 0.00 0.00 0.00 0.00 0.00 IE FINER-ORD 78.00 9.58 0.18 2.23 18.00 0.00 0.00 28.30 8.30 0.04 0.00 0.00 0.00 5.35 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA Headlines 78.00 0.00 47.32 84.33 60.00 0.00 0.00 85.42 82.23 97.08 85.10 82.21 94.69 71.14 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA TSA 61.00 0.00 0.85 63.82 42.00 32.34 32.34 42.98 43.40 81.70 85.11 60.00 86.38 54.89 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA XBRL-Math 68.00 68.89 67.78 27.78 62.00 0.00 11.11 64.44 44.44 7.78 3.33 2.22 6.67 14.44 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA FinQA 5.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 7.41 0.00 0.00 0.00 1.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA TATQA 0.00 0.00 0.00 0.36 44.00 0.00 0.00 0.05 1.73 4.14 0.00 0.00 0.00 15.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG ECTSUM 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG EDTSUM 25.00 19.13 16.80 16.59 18.00 0.98 0.10 20.16 23.89 19.92 12.49 8.06 2.06 13.61 0.00 0.00 0.00 0.00 0.00 0.00 0.00 RM CCF 52.50 50.00 50.62 51.34 50.00 50.93 50.00 52.94 50.31 50.05 50.00 50.00 51.18 50.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 PO BigData22 48.50 50.00 50.93 46.91 50.00 50.75 50.00 49.89 51.82 50.80 53.12 50.00 52.12 50.26 0.00 0.00 0.00 0.00 0.00 0.00 0.00 DM MSFT 41.32 65.06 0.00 0.00 72.25 74.03 79.97 49.32 0.00 0.00 68.81 74.50 66.53 65.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 ZH IE RRE 63.25 0.00 67.52 54.70 46.15 36.75 36.75 8.55 7.69 0.85 2.56 0.85 0.85 2.56 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA AIE 82.26 0.00 82.01 80.99 76.80 33.82 33.82 83.03 80.17 40.81 10.04 4.32 21.55 54.48 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA LNE 63.30 0.00 58.72 55.50 41.28 9.17 9.17 57.80 59.17 29.82 22.48 12.84 32.11 26.61 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA FinancialIQ 32.53 0.00 35.52 66.83 62.71 25.19 25.20 77.09 65.32 26.21 57.07 34.70 31.48 40.52 0.00 0.00 0.00 0.00 0.00 0.00 0.00 JA TA chabsa 0.00 0.00 0.00 48.43 32.17 8.98 23.96 4.54 44.35 46.94 47.59 23.96 57.36 34.62 0.00 0.00 0.00 0.00 0.00 0.00 0.00 ES TA MultiFin 61.74 0.00 53.91 62.17 48.26 22.17 22.17 46.52 46.96 43.04 31.74 12.61 44.78 51.30 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA TSA 0.39 0.00 29.17 52.29 24.29 63.04 63.46 31.63 46.46 31.03 68.19 63.38 16.64 51.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA EFPA 31.14 0.00 18.86 67.54 66.67 25.44 25.44 65.79 55.70 32.46 65.79 25.44 91.67 48.25 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG FNS-2023 25.94 18.11 0.00 9.61 12.14 0.00 0.00 5.93 7.50 1.64 5.71 10.62 1.65 9.27 0.00 0.00 0.00 0.00 0.00 0.00 0.00 EL IE GRFinNUM 9.18 20.98 7.43 49.12 46.34 0.00 0.00 36.77 0.40 0.00 0.00 0.00 0.00 70.06 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA GRFinQA 78.22 0.00 50.00 74.22 64.44 22.67 22.67 60.44 48.89 25.33 57.78 28.44 23.11 64.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TG GRFNS-2023 25.50 16.95 37.72 16.90 13.61 0.24 0.21 9.71 5.60 11.20 6.48 14.45 3.56 34.46 0.00 0.00 0.00 0.00 0.00 0.00 0.00 TA GRMultiFin 59.26 0.00 61.11 55.56 50.00 38.89 38.89 70.37 38.89 35.19 53.70 40.74 35.19 72.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 BI TG DOLFIN 92.29 90.13 86.26 89.17 92.13 35.92 35.92 92.29 91.80 69.24 91.60 71.81 66.57 91.59 0.00 0.00 0.00 0.00 0.00 0.00 0.00 MU QA PolyFiQA-Easy 9.79 9.56 34.72 27.73 25.04 15.02 14.74 19.34 18.81 2.44 2.40 11.63 0.63 7.06 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA PolyFiQA-Expert 5.31 4.85 30.35 20.60 18.56 13.83 16.01 18.17 16.35 6.38 0.71 8.80 0.00 9.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 40.98 14.59 30.61 39.50 42.20 19.34 20.41 38.07 33.06 26.83 31.24 24.40 28.95 35.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Vision EN IE EnglishOCR 21.38 0.00 0.00 12.39 0.00 10.70 11.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 QA TableBench 66.70 0.00 0.00 32.30 0.00 28.60 60.90 0.00 74.90 0.00 0.00 0.00 0.00 0.00 59.30 57.30 0.00 0.00 0.00 0.00 0.00 JA IE JapaneseOCR 21.63 0.00 0.00 24.52 0.00 25.82 26.72 0.00 21.59 0.00 0.00 0.00 9.70 6.62 0.00 0.00 0.00 0.00 0.00 0.00 0.00 ES IE SpanishOCR 78.55 0.00 0.00 4.12 0.00 5.60 4.40 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 EL IE GreekOCR 41.86 0.00 0.00 42.51 0.00 23.69 30.60 0.00 0.00 0.00 0.00 0.00 8.25 6.80 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 46.02 0.00 0.00 23.17 0.00 18.88 26.80 0.00 19.30 0.00 0.00 0.00 0.00 0.00 15.45 14.14 0.00 0.00 0.00 0.00 0.00 Audio EN TG MDRM-test 95.77 0.00 0.00 0.00 0.00 0.00 0.00 0.00 96.43 0.00 0.00 0.00 0.00 0.00 0.00 0.00 97.86 96.03 95.32 48.48 49.17 TG FinAudioSum 6.30 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.30 0.00 4.80 0.00 0.00 Average 51.04 0.00 0.00 0.00 0.00 0.00 0.00 0.00 48.22 0.00 0.00 0.00 0.00 0.00 0.00 0.00 51.58 48.02 50.06 24.24 24.59 Modality-Balanced Average 46.01 4.86 10.20 20.89 14.07 12.74 15.74 12.69 33.53 8.94 10.41 8.13 9.65 11.84 5.15 4.71 17.19 16.01 16.69 8.08 8.20
Table 3: Standardized performance of evaluated LLMs on the 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. The final row presents the modality-balanced average, computed as the mean of text, vision, and audio averages.
Figure 3: Performance across modalities: Text, Vision, Audio.
Figure 4: Performance across languages: EN, ZH, JA, ES, EL, BI, MU.